OpenStudy (idealist10):

Find all solutions of y'-xy=x^3*y^3.

OpenStudy (idealist10):

v=y^-2=1/y^2 y=1/sqrt(v) -1/(2v^(3/2))*dv/dx-x/sqrt(v)=x^3*1/(v*sqrt(v)) dv/dx+2xv=-2x^3 u(x)=e^x^2 e^x^2*v'+(e^x^2)2xv=-2x^3*e^x^2

OpenStudy (idealist10):

@amistre64

OpenStudy (idealist10):

@thomaster

OpenStudy (idealist10):

@SolomonZelman @ash2326 @Ashleyisakitty @Luigi0210 @ganeshie8

OpenStudy (ash2326):

This is a linear differential equation, sorry I have forgotten this :(

OpenStudy (idealist10):

That's fine.

OpenStudy (ash2326):

Perhaps @amistre64 Can help.

OpenStudy (amistre64):

been long removed from me too with little to no practice

OpenStudy (amistre64):

all i can possible recollect is to try a reduction formula or seperate

OpenStudy (idealist10):

I used the method to solve a Bernoulli equation in the work that I provided above.

OpenStudy (amistre64):

burnnoodle is a reduction process yes

OpenStudy (amistre64):

there is some y/x method in the bac of my head but i hahvent a clue if we could apply it here

OpenStudy (amistre64):

y'-xy=x^3*y^3 y = vx, y' = v+v'x v +v'x -x(vx)=x^3*(vx)^3 v +v'x -vx^2=v^3 x^6 prolly not

OpenStudy (idealist10):

I have never seen that before.

OpenStudy (amistre64):

ive prolly never seen it before either, hard to tell with this memory

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (johnweldon1993):

Ahh wait...I do believe you are on the ight track..now my memory is shaky as well but from the: \[\large (e^x)^2 \times \frac{dv}{dx}+((e^x)^2)2xv=-2x^3(e^x)^2\] The reverse product rule here would give \[\large \frac{d}{dx}((e^x)^2v(x)) = -2(e^x)^2x^3\] And integrate both sides to get \[\large (e^x)^2v(x) = -(e^x)^2(x^2 + 1)\]

OpenStudy (johnweldon1993):

sorry add in the \(\large + c_1\) at the end

OpenStudy (idealist10):

That's where I was stucked, how do I integrate -2x^3*e^x^2? I used integration by parts but it only complicated the problem more.

OpenStudy (amistre64):

we have a polly that goes to 0, and an exponential that integrates if thats the case

OpenStudy (amistre64):

pfft, e^x^2 ... is that ^(x^2)? or (^x)^2

OpenStudy (idealist10):

(-2x^3)(e^(x^2))----> I need to integrate this.

OpenStudy (johnweldon1993):

I did it as 

OpenStudy (amistre64):

e^(x^2) does have a taylor polynomial representation

OpenStudy (johnweldon1993):

And I integrated using a u-sub first then by parts \[\large \int -2(e^{x^2})x^3\] Let u = x^2 du = 2xdx \[\large - \int e^u udu\] *udu because 2xdx * x^2 = 2x^3dx so we have the x^3 taken care of... now we can do it by parts

OpenStudy (johnweldon1993):

Letting s = u ds = du v = e^u dv = e^udu So we would have \[\large -ue^u +\int e^udu\] And we know the integral there so we would have \[\large e^u - e^uu + c\] plug back in x^2 for 'u' and we have \[\large e^{x^2} - e^{x^2}x^2 + c\] Factor out a -e^(x^2) \[\large -e^{x^2}(x^2 - 1) + c\]

OpenStudy (johnweldon1993):

god that was a hassle to type >.< lol sorry for the wait

OpenStudy (idealist10):

Thanks!